OpenStudy (anonymous):

identify the maximum and minimum values of the function y = 3 cos x in the interval [-2π, 2π]. Use your understanding of transformations, not your graphing calculator.

OpenStudy (zehanz):

Do you know the min and max values of just cos x on that interval?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i am just completely lost on this one

OpenStudy (yacoub1993):

the largest cosine can be is 1 and the smallest it can be is -1

OpenStudy (yacoub1993):

and so what is the largest 3cos(x) can be?

OpenStudy (anonymous):

3 ?

OpenStudy (anonymous):

no my answer will be 3 and -3 ?

OpenStudy (anonymous):

The minimum of f(x) = cos(x) is when x = pi + k*2pi where k is an integer, and this minimum is -1, and the maximum of f(x) = cos(x) is when x = k*2pi, and this maximum is +1. Multiply f(x) by 3 and you get a new function g(x) = 3*cos(x), so the minimum and maximum values of g(x) are now 3*-1 and 3*1, respectively.

OpenStudy (anonymous):

oh ok i think i understand

OpenStudy (anonymous):

Yey!

OpenStudy (anonymous):

thank you :)